OpenStudy (anonymous):

Help me solve for x plz =D \[0=89900\left(\frac1{x^2}+\frac1{(8.0-x)^2}\right)\]

OpenStudy (anonymous):

\[\frac{89900}{x^2}=-\frac{89900}{(8.0-x)^2}\]

OpenStudy (anonymous):

\[1=\frac{x^2}{(8.0-x)^2}\]

zepdrix (zepdrix):

\[\large 0=89900\left(\frac1{x^2}+\frac1{(8.0-x)^2}\right)\]Getting a common denominator in the brackets gives us,\[\large 0=89900\left(\frac{(8.0-x)^2+x^2}{x^2(8.0-x)^2}\right)\]Divide both sides by 89900, set the numerator equal to zero to solve for x.\[\large 0=(8.0-x)^2+x^2\] It looks like you're headed in the same direction with the way you approached it though :D From here we just have a quadratic.

zepdrix (zepdrix):

Which doesn't appear to have any real solutions :\ hmm

OpenStudy (anonymous):

makes sense. Thanks....my algebra sucks :S just had to double check so from here we have \[8^2-16x+x^2+x^2\] \[=2x^2 -16x+64\] \[=2(x^2-8x+32)\] The answer should be 4

zepdrix (zepdrix):

I put it into Wolfram just to check. It's giving me the same answer I came up with. Throwing this into the `Quadratic Formula` should give, \[\large x=4 \pm 4i\] Yer saying it's suppose to be 4 though? :o

OpenStudy (anonymous):

yep

zepdrix (zepdrix):

Hmm, yah it's not 4 I'm afraid :\ Here is the problem worked out on wolfram if that helps. http://www.wolframalpha.com/input/?i=0%3D89900%281%2Fx%5E2%2B1%2F%288-x%29%5E2%29

OpenStudy (anonymous):

oh well...I'm too tired to care...haha

zepdrix (zepdrix):

XD